OpenStudy (anamishka):

A ramp is used to move crates of apples to a loading dock at different heights. When the horizontal distance AB is 12 feet, the height of the loading dock, BC, is 6 feet. What is the height of loading dock DE? (AB=12 BD=22 ABD=34 BC=6 DE=?) A. 10 feet B. 12 feet C. 17 feet D. 18 feet

OpenStudy (retireed):



OpenStudy (anamishka):

Thank you! I understand it now! The Answer is C 17 feet. X/34 = 6/12 X=17

OpenStudy (retireed):

Yes it was very difficult to figure out what the question was trying to explain.

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)